Strontium-90 is a dangerous by-product of atomic testing because it mimics the action of calcium in the body. It decay in two B-emissions to give zirconium-90 (nuclear mass = 89.8824 g). a) write a balanced nuclear equation for the overall decat of Sr-90. b) calculate ∆m in grams when one mole of Sr-90 decays to Zr-90.
c) How much energy (in KJ) is given off by the decay of 6.50 mg of Sr-90?
1
Expert's answer
2020-11-14T13:52:23-0500

a) Sr(38) = Zr(40) + 2B-

b) 9.11 x 10-31 x 2 = 1,822e-27 g.

c) e=mc^2 = 0,0065 x (3x10^8)^2 = 5,85e11 kJ